Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 2/5/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
    • Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • This needs to be cross-checked - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it in situ before the run?
      • Josef :"we just run 50A through it when it is connected to the magnet"
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
    • CWhyte to contact DL to clarify their Hall Task List items
    • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
      • Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC
    • HNebrensky to send BLOC-chasing email
  • DC Update
  • LH2 Update
  • Tracker Update
  • List of Hall activities in the shutdown (Colin)
  • Schedule for startup
    • Startup run on Monday May 8th - Hall should be reserved for operations activities all day. Controlled Entry is possible
      but only when permitted by the MOM.
    • We will not be running the SS magnets this user cycle, or taking data with Neon.
    • We will need to run the FC at some point.
    • Co-ordination between DC and MOM will be key.

AOB:

Minutes

  • Actions
    • Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • This needs to be cross-checked - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it in situ before the run?
      • Josef :"we just run 50A through it when it is connected to the magnet"
      • Run into short in next couple of days
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
    • CWhyte to contact DL to clarify their Hall Task List items
    • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
      • Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC
      • Still need to decide, then get Trevor to re-connect if needed
    • HNebrensky to send BLOC-chasing email
  • DC Update
    Tracker cryo lost channels - possibly water ingress, looking for sample bottles.
    DC to check that CC magnet supplies are all into shorts.
  • LH2 Update
    HAZOP on 16th/17th May; have the panel; hope to get LOPA done at end.
    Working for the Neon test - will be "as LH2". Several days of cooling, and 1 day with liquid Neon. Quench test after run.
    The LH2 Beacon will be live during the Neon test - should warn ISIS MCR beforehand.
  • Tracker Update
    One of the boards needs changing - Ed will do while he's down.
    • Tracker volume Helium purge: MU has added more putty round the leaky one.
      When should we start the flow? Should estimate from volume volume and flowrate how long it takes to change the gas in the Tracker, then allow 5 volume changes before running.
      Craig to show MOM how to start flow and change bottles
  • Schedule for startup
    • Startup run on Monday May 8th - Hall should be reserved for operations activities all day. Controlled Entry is possible
      but only when permitted by the MOM.
    • We will not be running the CC magnets this user cycle, or taking data with Neon.
    • The PRY will stay open.
    • Co-ordination between DC and MOM will be key.
      • Run meetings 10 a.m. in ISIS coffee lounge
  • AOB
    CRogers to check the M3 currents page and send to MOM. Also the CC current setting sign-off, even if it is null.
    CWhyte to request geometry survey of SS's and FC, and check TOF1 status.
    DS restart later this week
    EO hopes to have target1ctl finished by end Friday - get Joe down week after next.
    Shifter training should always include some level of Controlled Entry refresher.
    MOM is about to send off the Run Request to ISIS.

Updated by Nebrensky, Henry over 6 years ago · 12 revisions